OpenStudy (anonymous):

Mike bungee jumped from a building with eight stories each story 15 and 4/5 feet in height how tall was building?

OpenStudy (anonymous):

15 4/5 can be written as 79/5(15*5=75+4=79) so each storey is 79/5 feet high..so height of 8 stories would be...(79/5)*8

OpenStudy (anonymous):

@deewash can you calculate now

OpenStudy (anonymous):

@deewash ..are you working it out..do you need my help?

OpenStudy (anonymous):

Need help